OpenStudy (wesdg1978):

For x^2-y^2=5, dy/dx=x/y How do I then find d^2y/dx^2?

OpenStudy (anonymous):

that is the second derivitive..

OpenStudy (wesdg1978):

For the second derivative, I get (y-x)/y^2 , but that's not an answer choice.

OpenStudy (anonymous):

what are the choices

OpenStudy (wesdg1978):

Nor is the answer you posted.

OpenStudy (anonymous):

what are the choices

OpenStudy (wesdg1978):

A) dy/dx =x/y; d2y/dx2 =(y - x2)/y2 B) dy/dx =x/y; d2y/dx2 =(y2 - x2)/y3 C) dy/dx =x/y; d2y/dx2 =(y2 - x2)/y2 D) dy/dx = -x/y; d2y/dx2 =(y2 - x2)/y3

OpenStudy (anonymous):

oh use quoitent rule

OpenStudy (wesdg1978):

That gets (y-x)/y^2

OpenStudy (anonymous):